OpenStudy (yourboydusty):

What is the slope of the line given by the equation below? y + 5 = -22(x + 2) How do i solve the equation, if you can show the steps of the problem it would be very helpful. Thank You!

OpenStudy (s4sensitiveandshy):

\[y-y_1 =m(x -x_1)\] the given equation is in point slope form where m is slope \[(x_1 ,y_1)\] given point

OpenStudy (yourboydusty):

im still kinda confused..sorry

OpenStudy (s4sensitiveandshy):

compare the given equation with the one i posted above^^ what number is replaced by m ?

OpenStudy (yourboydusty):

-22

OpenStudy (s4sensitiveandshy):

yes that's correct that's the slope of that equation.

OpenStudy (yourboydusty):

but how do i actually solve it

OpenStudy (s4sensitiveandshy):

solve for what ? do you want to rewrite it in slope intercept form ? which is y=mx+b

OpenStudy (yourboydusty):

oh...wait i miss-read the question...sorry...thank you for the help :)

OpenStudy (s4sensitiveandshy):

np :) you don't have to solve for anything to find the slope from point slope form equation